How to simplify (1 + cos theta) (1 - cos theta) ?
Answer by John10(297)

How to simplify (1 + cos theta) (1 - cos theta) ?
Use the formula a^2 - b^2 = (a-b)(a+b)
(1 + cos theta) (1 - cos theta)
= (1)^2 - (cos theta)^2
= 1 - (cos theta)^2
= (sin theta)^2 [because (sin x)^2 + (cos x)^2 = 1]
Hope it helps!John10